# Scalable MatMul-free Language Modeling

### v5

@@ 3.2 Hardware-efficient Fused BitLinear Layer
BitNet showed that stabilizing ternary layers requires an additional RMSNorm before the BitLinear input (for more details, refer to Appendix A). However, the vanilla implementation of BitNet is not efficient. Modern GPUs feature a memory hierarchy with a large, global high-bandwidth memory (HBM) and smaller, faster shared memory (SRAM), and the implementation of BitNet introduced many I/O operations: reading the previous activation into SRAM for RMSNorm, writing it back for quantization, reading it again for quantization, storing it, and reading it once more for the Linear operation. To address this inefficiency, we read the activation only once and apply RMSNorm and quantization as fused operations in SRAM.
Algorithm 1 presents our approach for improving the hardware efficiency of the BitLinear layer by fusing quantized RMSNorm and BitLinear operations. Optimal utilization of SRAM to reduce HBM I/O costs can significantly speed up computations. Since the activations in this model have a larger memory footprint than weights due to ternary weights and the amount of element-wise operations, our optimization efforts focus on activations.
The forward_pass function in Algorithm 1 outlines the forward pass of our fused BitLinear layer. It first calls rms_norm_fwd to perform RMSNorm on input activations $\mathbf{X}$, loading normalized activations $\mathbf{Y}$, mean $\mu$, variance $\sigma^{2}$, and scaling factor $r$ from HBM. The normalized activations $\mathbf{Y}$ are then quantized to obtain $\widetilde{\mathbf{Y}}$ and the weights $\mathbf{W}$ are quantized using weight_quant, with both performed without off-chip data movement. Finally, the output $\mathbf{O}$ is computed on-chip by multiplying quantized activations $\widetilde{\mathbf{Y}}$ with the ternary weights $\widetilde{\mathbf{W}}$, adding the bias $\boldsymbol{b}$, and then storing the result back to HBM.
Figure 2: Overview of the MatMul-free LM. The sequence of operations are shown for vanilla self-attention (top-left), the MatMul-free token mixer (top-right), and Ternary Accumulations. The MatMul-free LM employs a MatMul-free token mixer (MLGRU) and a MatMul-free channel mixer (MatMul-free GLU) to maintain the transformer-like architecture while reducing compute cost.
The backward_pass function first loads $\mathbf{X}$, $\mathbf{W}$, $\boldsymbol{b}$, $\mathbf{O}$, $\mu$, $\sigma^{2}$, $r$, and the output gradient $\mathrm{d}\mathbf{O}$ from HBM. The gradient $\mathrm{d}\mathbf{Y}$ is then computed on-chip by multiplying the output gradient $\mathrm{d}\mathbf{O}$ with the transposed weight matrix $\mathbf{W}^{\top}$. Next, it calls rms_norm_bwd on-chip to backpropagate through RMSNorm, computing the input gradient $\mathrm{d}\mathbf{X}$. The weight gradient $\mathrm{d}\mathbf{W}$ is calculated on-chip by multiplying the transposed output gradient $\mathrm{d}\mathbf{O}^{\top}$ with the quantized activations $\widetilde{\mathbf{Y}}$, and the bias gradient $\mathrm{d}\boldsymbol{b}$ is obtained by summing $\mathrm{d}\mathbf{O}$. The computed gradients $\mathbf{dX}$, $\mathrm{d}\mathbf{W}$, and $\mathrm{d}\boldsymbol{b}$ are then stored back to HBM. Sec. 4.4 presents an experimental comparison between Vanilla BitLinear and Fused BitLinear.
@@ 3.3 MatMul-free Language Model Architecture
We adopt the perspective from Metaformer [27], which suggests that Transformers consist of a token-mixer (for mixing temporal information, i.e., Self Attention [28], Mamba [29]) and a channel-mixer (for mixing embedding/spatial information, i.e., feed-forward network, Gated Linear Unit (GLU)  [30, 31]). A high-level overview of the architecture is shown in Fig. 2.
@@ 3.3.1 MatMul-free Token Mixer
Self-attention is the most common token mixer in modern language models, relying on matrix multiplication between three matrices: $Q$, $K$, and $V$. To convert these operations into additions, we binarize or ternarize at least two of the matrices. Assuming all dense layer weights are ternary, we quantize $Q$ and $K$, resulting in a ternary attention map that eliminates multiplications in self-attention. However, as shown in Fig. 1, the model trained this way fails to converge. One possible explanation is that activations contain outliers crucial for performance but difficult to quantize effectively [32, 33]. To address this challenge, we explore alternative methods for mixing tokens without relying on matrix multiplications.
By resorting to the use of ternary RNNs, which combine element-wise operations and accumulation, it becomes possible to construct a MatMul-free token mixer. Among various RNN architectures, the GRU is noted for its simplicity and efficiency, achieving similar performance to Long Short-Term Memory (LSTM) [34] cells while using fewer gates and having a simpler structure. Thus, we choose the GRU as the foundation for building a MatMul-free token mixer. We first revisit the standard GRU and then demonstrate, step by step, how we derive the MLGRU.
@@
The GRU [13] is a widely-used variant of the RNN architecture that is simpler and more computationally efficient compared to the LSTM unit while still maintaining comparable performance. The GRU can be formalized as follows:
$\boldsymbol{r}_{t}$ $=\sigma\left(\boldsymbol{x}_{t}\mathbf{W}_{xr}+\boldsymbol{h}_{t-1}\mathbf{W}_{hr}+\mathbf{b}_{r}\right)\in\mathbb{R}^{1\times d},$ (1)
$\boldsymbol{f}_{t}$ $=\sigma\left(\boldsymbol{x}_{t}\mathbf{W}_{xf}+\boldsymbol{h}_{t-1}\mathbf{W}_{hf}+\mathbf{b}_{f}\right)\in\mathbb{R}^{1\times d},$ (2)
$\boldsymbol{c}_{t}$ $=\tanh\left(\boldsymbol{x}_{t}\mathbf{W}_{xc}+(\boldsymbol{r}_{t}\odot\boldsymbol{h}_{t-1})\mathbf{W}_{cc}+\mathbf{b}_{c}\right)\in\mathbb{R}^{1\times d},$ (3)
$\boldsymbol{h}_{t}$ $=\boldsymbol{f}_{t}\odot\boldsymbol{h}_{t-1}+(1-\boldsymbol{f}_{t})\odot\boldsymbol{c}_{t}\in\mathbb{R}^{1\times d},$ (4)
@@
where $\boldsymbol{x}_{t}\in\mathbb{R}^{1\times m}$ is the input vector at time step $t$, $\boldsymbol{h}_{t-1}\in\mathbb{R}^{1\times d}$ is the hidden state vector from the previous time step, $\boldsymbol{r}_{t}\in\mathbb{R}^{1\times d}$ is the reset gate vector, $\boldsymbol{f}_{t}\in\mathbb{R}^{1\times d}$ is the forget gate vector, $\boldsymbol{c}_{t}\in\mathbb{R}^{1\times d}$ is the candidate hidden state, $\boldsymbol{h}_{t}\in\mathbb{R}^{1\times d}$ is the final hidden state vector at time step $t$, $\boldsymbol{o}_{t}\in\mathbb{R}^{1\times d}$ is the output vector at time step $t$, $\mathbf{W}{(\cdot)}\in\mathbb{R}^{m\times d}$ and $\mathbf{b}{(\cdot)}\in\mathbb{R}^{1\times d}$ are learnable weight matrices and bias vectors, respectively, $\sigma(\cdot)$ is the sigmoid activation function, and $\odot$ denotes element-wise multiplication.
A key characteristic of the GRU is the coupling of the input gate vector $\boldsymbol{f}_{t}$ and the forget gate vector $(1-\boldsymbol{f}_{t})$, which together constitute the ‘leakage’ unit. This leakage unit decays the hidden state $\boldsymbol{h}_{t-1}$ and the candidate hidden state $\boldsymbol{c}_{t}$ through element-wise multiplication, see Eq. 4. This operation allows the model to adaptively retain information from the previous hidden state $\boldsymbol{h}_{t-1}$ and incorporate new information from the candidate hidden state $\boldsymbol{c}_{t}$. Importantly, this operation relies solely on element-wise multiplication, avoiding the need for the MatMul. We aim to preserve this property of the GRU while introducing further modifications to create a MatMul-free variant of the model.
@@
We first remove hidden-state related weights $\mathbf{W}_{cc}$, $\mathbf{W}_{hr}$, $\mathbf{W}_{hf}$, and the activation between hidden states ($\tanh$). This modification not only makes the model MatMul-free but also enables parallel computation similar to Transformers. This approach is critical for improving computational efficiency, as transcendental functions are expensive to compute accurately, and non-diagonal matrices in the hidden-state would hinder parallel computations. This modification is a key feature of recent RNNs, such as the Linear Recurrent Unit [35], Hawk [36], and RWKV-4 [37]. We then add a data-dependent output gate between $\boldsymbol{h}_{t}$ and $\boldsymbol{o}_{t}$, inspired by the LSTM and widely adopted by recent RNN models:
$\boldsymbol{g}_{t}$ $=\sigma(\boldsymbol{x}_{t}\mathbf{W}_{g}+\mathbf{b}_{g})\in\mathbb{R}^{1\times d},$
$\boldsymbol{o}^{\prime}_{t}$ $=\boldsymbol{g}_{t}\odot\boldsymbol{h}_{t}\in\mathbb{R}^{1\times d},$
$\boldsymbol{o}_{t}$ $=\boldsymbol{o}^{\prime}_{t}\mathbf{W}_{o}+\mathbf{b}_{o}\in\mathbb{R}^{1\times d}.$
@@
Following the approach of HGRN [38], we further simplify the computation of the candidate hidden state by keeping it as a simple linear transform, rather than coupling it with the hidden state. This can be rewritten as a linear transformation of the input. Finally, we replace all remaining weight matrices with ternary weight matrices, completely removing MatMul operations. The resulting MLGRU architecture can be formalized as follows:
$\boldsymbol{f}_{t}$ $=\sigma\left(\boldsymbol{x}_{t}\circledast\mathbf{W}_{f}+\mathbf{b}_{f}\right)\in\mathbb{R}^{1\times d},$
$\boldsymbol{c}_{t}$ $=\tau\left(\boldsymbol{x}_{t}\circledast\mathbf{W}_{c}+\mathbf{b}_{c}\right)\in\mathbb{R}^{1\times d},$
$\boldsymbol{h}_{t}$ $=\boldsymbol{f}_{t}\odot\boldsymbol{h}_{t-1}+(1-\boldsymbol{f}_{t})\odot\boldsymbol{c}_{t}\in\mathbb{R}^{1\times d},$
$\boldsymbol{g}_{t}$ $=\sigma(\boldsymbol{x}_{t}\circledast\mathbf{W}_{g}+\mathbf{b}_{g})\in\mathbb{R}^{1\times d},$
$\boldsymbol{o}^{\prime}_{t}$ $=\boldsymbol{g}_{t}\odot\boldsymbol{h}_{t}\in\mathbb{R}^{1\times d},$
$\boldsymbol{o}_{t}$ $=\boldsymbol{o}^{\prime}_{t}\circledast\mathbf{W}_{o}+\mathbf{b}_{o}\in\mathbb{R}^{1\times d}.$
@@
where $\mathbf{W}_{c},\mathbf{W}_{f},\mathbf{W}_{o},\mathbf{W}_{g}\in\mathbb{R}^{d\times d}$ are ternary weights, $\boldsymbol{f}_{t}$ is the forget gate output, $\sigma$ is the Sigmoid activation function, $\boldsymbol{c}_{t}$ is the input vector, $\tau$ is the SiLU activation function, $\boldsymbol{h}_{t}$ is the hidden state, $\boldsymbol{g}_{t}$ is the output gate, $\boldsymbol{o}^{\prime}_{t}$ is the intermediate output, and $\boldsymbol{o}_{t}$ is the final output at time step $t$. The initial hidden state $\boldsymbol{h}_{0}$ is set to $\mathbf{0}$. Similarly to HGRN, we also employ the $\mathrm{cummax}$ operation to bound the forget gate values in deeper layers closer to 1, though omit this above for brevity. The MLGRU can be viewed as a simplified variant of HGRN that omits complex-valued components and reduces the hidden state dimension from $2d$ to $d$. This simplification makes MLGRU more computationally efficient while preserving essential gating mechanisms and ternary weight quantization.
Alternatively to the MLGRU, which employs a data-dependent decay with element-wise product hidden state, the a similarly modified version of the RWKV-4 model can also satisfy the requirement of a MatMul-free token mixer, utilizing static decay and normalization. The performance of using RWKV-4 as a MatMul-free token mixer is discussed in the Experiment section, with a detailed description of the RWKV-4 model provided in Appendix B. However, RWKV-4 introduces exponential and division operations, which are less hardware-efficient compared to the MLGRU.
@@ 3.3.2 MatMul-free Channel Mixer
For MatMul-free channel mixing, we use GLU, which is widely adopted in many modern LLMs, including Llama [39, 40, 41], Mistral [42] and RWKV [37], and a BitLinear-adapted version can be expressed as follows:
$\boldsymbol{g}_{t}$ $=\boldsymbol{x}_{t}\circledast\boldsymbol{W}_{g}\in\mathbb{R}^{1\times l},$
$\boldsymbol{u}_{t}$ $=\boldsymbol{x}_{t}\circledast\mathbf{W}_{u}\in\mathbb{R}^{1\times l},$
$\boldsymbol{p}_{t}$ $=\tau(\boldsymbol{g}_{t})\odot\boldsymbol{u}_{t}\in\mathbb{R}^{1\times l},$
$\boldsymbol{d}_{t}$ $=\boldsymbol{p}_{t}\circledast\mathbf{W}_{d}\in\mathbb{R}^{1\times d},$
@@
The GLU consists of three main steps: 1) upscaling the $t$-step input $\boldsymbol{x}_{t}\in\mathbb{R}^{1\times d}$ to $\boldsymbol{g}_{t},\boldsymbol{u}_{t}\in\mathbb{R}^{1\times l}$ using weight matrices $\mathbf{W}_{g},\mathbf{W}_{u}\in\mathbb{R}^{d\times l}$ 2) elementwise gating $\boldsymbol{u}_{t}$ with $\boldsymbol{g}_{t}$ followed by a nonlinearity $f(\cdot)$, where we apply $\mathrm{Swish}$ [31]. 3) Down-scaling the gated representation $\boldsymbol{p}_{t}$ back to the original size through a linear transformation $\mathbf{W}_{d}$. Following Llama [39], we maintain the overall number of parameters of GLU at $8d^{2}$ by setting the upscaling factor to $\frac{8}{3}d$.

### v6

@@ 6.2 Hardware-efficient Fused BitLinear Layer
BitNet showed that stabilizing ternary layers requires an additional RMSNorm before the BitLinear input. However, the vanilla implementation of BitNet is not efficient. Modern GPUs feature a memory hierarchy with a large, global high-bandwidth memory (HBM) and smaller, faster shared memory (SRAM), and the implementation of BitNet introduced many I/O operations: reading the previous activation into SRAM for RMSNorm, writing it back for quantization, reading it again for quantization, storing it, and reading it once more for the Linear operation. To address this inefficiency, we read the activation only once and apply RMSNorm and quantization as fused operations in SRAM, which we present in Algorithm 1. Optimal utilization of SRAM to reduce HBM I/O costs can significantly speed up computations. Since the activations in this model have a larger memory footprint than ternary weights and the large number of element-wise operations, our optimization efforts focus on activations.
@@
The forward_pass function in Algorithm 1 first calls rms_norm_fwd to perform RMSNorm on input activations $\mathbf{X}$, loading normalized activations $\mathbf{Y}$, mean $\mu$, variance $\sigma^{2}$, and scaling factor $r$ from HBM. The normalized activations $\mathbf{Y}$ are then quantized to obtain $\widetilde{\mathbf{Y}}$ and the weights $\mathbf{W}$ are quantized using weight_quant, with both performed without off-chip data movement. Finally, the output $\mathbf{O}$ is computed on-chip by multiplying quantized activations $\widetilde{\mathbf{Y}}$ with the ternary weights $\widetilde{\mathbf{W}}$, adding the bias $\boldsymbol{b}$, and then storing the result back to HBM.
The backward_pass function first loads $\mathbf{X}$, $\mathbf{W}$, $\boldsymbol{b}$, $\mathbf{O}$, $\mu$, $\sigma^{2}$, $r$, and the output gradient $\mathrm{d}\mathbf{O}$ from HBM. The gradient $\mathrm{d}\mathbf{Y}$ is then computed on-chip by multiplying the output gradient $\mathrm{d}\mathbf{O}$ with the transposed weight matrix $\mathbf{W}^{\top}$. Next, it calls rms_norm_bwd on-chip to backpropagate through RMSNorm, computing the input gradient $\mathrm{d}\mathbf{X}$. The weight gradient $\mathrm{d}\mathbf{W}$ is calculated on-chip by multiplying the transposed output gradient $\mathrm{d}\mathbf{O}^{\top}$ with the quantized activations $\widetilde{\mathbf{Y}}$, and the bias gradient $\mathrm{d}\boldsymbol{b}$ is obtained by summing $\mathrm{d}\mathbf{O}$. The computed gradients $\mathbf{dX}$, $\mathrm{d}\mathbf{W}$, and $\mathrm{d}\boldsymbol{b}$ are then transferred back to HBM. Sec. 5.1 presents an experimental comparison between a vanilla BitLinear implementation and Fused BitLinear.
@@ 6.3 MatMul-free Language Model Architecture
We adopt the perspective from Metaformer yu2022metaformer , which suggests that Transformers consist of a token-mixer (for mixing temporal information, i.e., Self Attention vaswani2017attention , Mamba gu2023mamba ) and a channel-mixer (for mixing embedding/spatial information, i.e., feed-forward network, Gated Linear Unit (GLU)  dauphin2017glu ; shazeer2020glu ). A high-level overview of the architecture is shown in Fig. 1.
@@ 6.3.1 MatMul-free Token Mixer
Self-attention is the most common token mixer in modern language models, relying on matrix multiplication between three matrices: $Q$, $K$, and $V$. To convert these operations into additions, we ternarize at least two of the matrices. Assuming all dense layer weights are ternary, we quantize $Q$ and $K$, resulting in a ternary attention map that eliminates multiplications in self-attention. However, as shown in Fig. 4, such a model fails to converge. One possible explanation is that activations contain outliers crucial for performance but difficult to quantize effectively pan2023smoothquant+ ; xiao2023smoothquant . To address this challenge, we explore alternative methods for mixing tokens without relying on matrix multiplications.
By resorting to the use of ternary RNNs, which combine element-wise operations and accumulation, it becomes possible to construct a MatMul-free token mixer. Among various RNN architectures, the GRU is noted for its simplicity and efficiency, achieving similar performance to Long Short-Term Memory (LSTM) hochreiter1997long cells while using fewer gates and having a simpler structure. Thus, we choose the GRU as the foundation for building a MatMul-free token mixer. We first revisit the standard GRU and then demonstrate, step by step, how we derive the MLGRU.
@@
The GRU cho2014learning can be formalized as follows:
$\boldsymbol{r}_{t}$ $=\sigma\left(\boldsymbol{x}_{t}\mathbf{W}_{xr}+\boldsymbol{h}_{t-1}\mathbf{W}_{hr}+\mathbf{b}_{r}\right)\in\mathbb{R}^{d},$ (1)
$\boldsymbol{f}_{t}$ $=\sigma\left(\boldsymbol{x}_{t}\mathbf{W}_{xf}+\boldsymbol{h}_{t-1}\mathbf{W}_{hf}+\mathbf{b}_{f}\right)\in\mathbb{R}^{d},$ (2)
$\boldsymbol{c}_{t}$ $=\tanh\left(\boldsymbol{x}_{t}\mathbf{W}_{xc}+(\boldsymbol{r}_{t}\odot\boldsymbol{h}_{t-1})\mathbf{W}_{cc}+\mathbf{b}_{c}\right)\in\mathbb{R}^{d},$ (3)
$\boldsymbol{h}_{t}$ $=\boldsymbol{f}_{t}\odot\boldsymbol{h}_{t-1}+(1-\boldsymbol{f}_{t})\odot\boldsymbol{c}_{t}\in\mathbb{R}^{d},$ (4)
@@
where $\boldsymbol{x}_{t}\in\mathbb{R}^{m}$ is the input vector at time step $t$, $\boldsymbol{h}_{t-1}\in\mathbb{R}^{d}$ is the hidden state vector from the previous time step, $\boldsymbol{r}_{t}\in\mathbb{R}^{d}$ is the reset gate vector, $\boldsymbol{f}_{t}\in\mathbb{R}^{d}$ is the forget gate vector, $\boldsymbol{c}_{t}\in\mathbb{R}^{d}$ is the candidate hidden state, $\boldsymbol{h}_{t}\in\mathbb{R}^{d}$ is the final hidden state vector at time step $t$, $\boldsymbol{o}_{t}\in\mathbb{R}^{d}$ is the output vector at time step $t$, $\mathbf{W}{(\cdot)}\in\mathbb{R}^{m\times d}$ and $\mathbf{b}{(\cdot)}\in\mathbb{R}^{d}$ are learnable weight matrices and bias vectors, respectively, $\sigma(\cdot)$ is the sigmoid activation function, and $\odot$ denotes element-wise multiplication.
A key characteristic of the GRU is the coupling of the input gate vector $\boldsymbol{f}_{t}$ and the forget gate vector $(1-\boldsymbol{f}_{t})$, which together constitute the ‘leakage’ unit. This leakage unit decays the hidden state $\boldsymbol{h}_{t-1}$ and the candidate hidden state $\boldsymbol{c}_{t}$ through element-wise multiplication (Eq. 4). This operation allows the model to adaptively retain information from the previous hidden state $\boldsymbol{h}_{t-1}$ and incorporate new information from the candidate hidden state $\boldsymbol{c}_{t}$. Importantly, this operation relies solely on element-wise multiplication, avoiding the need for MatMul. We preserve this property of the GRU while introducing further modifications to create a MatMul-free variant of the model.
@@
We first remove hidden-state related weights $\mathbf{W}_{cc}$, $\mathbf{W}_{hr}$, $\mathbf{W}_{hf}$, and the activation between hidden states ($\tanh$). This modification not only makes the model MatMul-free but also linearized the GRU through time, enabling training via a parallel scan gu2023mamba ; voelker2019legendre . This approach is critical for improving computational efficiency, as transcendental functions are expensive to compute accurately, and non-diagonal transition matrices increase runtime complexity from $\mathcal{O(n)}$ to $\mathcal{O(n^{3})}$. This modification is a key feature of recent RNNs, such as the Linear Recurrent Unit orvieto2023lru , Hawk de2024griffin , and RWKV-4 rwkv4 . We then add a data-dependent output gate between $\boldsymbol{h}_{t}$ and $\boldsymbol{o}_{t}$, inspired by the LSTM and widely adopted by recent RNN models:
$\boldsymbol{g}_{t}$ $=\boldsymbol{x}_{t}\mathbf{W}_{g}+\mathbf{b}_{g}\in\mathbb{R}^{d},$
$\boldsymbol{o}^{\prime}_{t}$ $=\boldsymbol{g}_{t}\odot\sigma(\boldsymbol{h}_{t})\in\mathbb{R}^{d},$
$\boldsymbol{o}_{t}$ $=\boldsymbol{o}^{\prime}_{t}\mathbf{W}_{o}+\mathbf{b}_{o}\in\mathbb{R}^{d}.$
@@
Following the approach of HGRN qin2024hgrn , we further simplify the computation of the candidate hidden state by keeping it as a simple linear transform, rather than coupling it with the hidden state. This can be rewritten as a linear transformation of the input. Finally, we replace all remaining weight matrices with ternary weight matrices, completely removing the MatMul operations. The resulting MLGRU architecture can be formalized as follows:
$\boldsymbol{f}_{t}$ $=\sigma\left(\boldsymbol{x}_{t}\circledast\mathbf{W}_{f}+\mathbf{b}_{f}\right)\in\mathbb{R}^{d},$
$\boldsymbol{c}_{t}$ $=\tau\left(\boldsymbol{x}_{t}\circledast\mathbf{W}_{c}+\mathbf{b}_{c}\right)\in\mathbb{R}^{d},$
$\boldsymbol{h}_{t}$ $=\boldsymbol{f}_{t}\odot\boldsymbol{h}_{t-1}+(1-\boldsymbol{f}_{t})\odot\boldsymbol{c}_{t}\in\mathbb{R}^{d},$
$\boldsymbol{g}_{t}$ $=\boldsymbol{x}_{t}\circledast\mathbf{W}_{g}+\mathbf{b}_{g}\in\mathbb{R}^{d},$
$\boldsymbol{o}^{\prime}_{t}$ $=\boldsymbol{g}_{t}\odot\sigma(\boldsymbol{h}_{t})\in\mathbb{R}^{d},$
$\boldsymbol{o}_{t}$ $=\boldsymbol{o}^{\prime}_{t}\circledast\mathbf{W}_{o}+\mathbf{b}_{o}\in\mathbb{R}^{d}.$
@@
where $\mathbf{W}_{c},\mathbf{W}_{f},\mathbf{W}_{o},\mathbf{W}_{g}\in\mathbb{R}^{d\times d}$ consist of ternary weights, $\boldsymbol{f}_{t}$ is the forget gate output, $\sigma$ is the Sigmoid activation function, $\boldsymbol{c}_{t}$ is the input vector, $\tau$ is the SiLU activation function, $\boldsymbol{h}_{t}$ is the hidden state, $\boldsymbol{g}_{t}$ is the output gate, $\boldsymbol{o}^{\prime}_{t}$ is the intermediate output, and $\boldsymbol{o}_{t}$ is the final output at time step $t$. The initial hidden state $\boldsymbol{h}_{0}$ is set to $\mathbf{0}$. Similarly to HGRN, we also employ the $\mathrm{cummax}$ operation to bound the forget gate values in deeper layers closer to 1, though omit this above for brevity. The MLGRU can be viewed as a simplified variant of HGRN that omits complex-valued components and reduces the hidden state dimension from $2d$ to $d$. This simplification makes MLGRU more computationally efficient while preserving essential gating mechanisms and ternary weight quantization.
Recurrent LLMs are known to have limitations in long-context benchmarks and retrieval. Linear recurrent models partially address this by 1) avoiding non-linearities through time to improve gradient flow, 2) introducing data-dependent decay, and 3) introducing lower-bounds on the ‘forget gate’ qin2024hgrn . MatMul-free LM adopts these best practices, and recent work shows that hybrid architectures with very few Transformer blocks can compensate for RNN performance across longer contexts wang2025systematic .
@@ 6.3.2 MatMul-free Channel Mixer
For MatMul-free channel mixing, we use GLU, which is widely adopted in many modern LLMs, including Llama touvron2023llama ; touvron2023llama2 ; llama3modelcard , Mistral jiang2023mistral and RWKV rwkv4 , and a BitLinear-adapted version can be expressed as follows:
$\boldsymbol{g}_{t}$ $=\boldsymbol{x}_{t}\circledast\boldsymbol{W}_{g}\in\mathbb{R}^{l},$
$\boldsymbol{u}_{t}$ $=\boldsymbol{x}_{t}\circledast\mathbf{W}_{u}\in\mathbb{R}^{l},$
$\boldsymbol{p}_{t}$ $=\tau(\boldsymbol{g}_{t})\odot\boldsymbol{u}_{t}\in\mathbb{R}^{l},$
$\boldsymbol{d}_{t}$ $=\boldsymbol{p}_{t}\circledast\mathbf{W}_{d}\in\mathbb{R}^{d},$
@@
The GLU consists of three main steps: 1) upscaling the $t$-step input $\boldsymbol{x}_{t}\in\mathbb{R}^{d}$ to $\boldsymbol{g}_{t},\boldsymbol{u}_{t}\in\mathbb{R}^{l}$ using weight matrices $\mathbf{W}_{g},\mathbf{W}_{u}\in\mathbb{R}^{d\times l}$ 2) elementwise gating $\boldsymbol{u}_{t}$ with $\boldsymbol{g}_{t}$ followed by a nonlinearity $f(\cdot)$, where we apply $\mathrm{Swish}$ shazeer2020glu . 3) Down-scaling the gated representation $\boldsymbol{p}_{t}$ back to the original size through a linear transformation $\mathbf{W}_{d}$. Following Llama touvron2023llama , we maintain the overall number of parameters of GLU at $8d^{2}$ by setting the upscaling factor to $\frac{8}{3}d$.